# The Effect of the Gotthard Base Tunnel on Road Traffic: A Synthetic Control Approach

### v1

@@ 6.1 Main Results
Figure 4 shows the main result when applying the synthetic control method. The trajectories of the Gotthard motorway section and the synthetic Gotthard motorway section mainly track each other closely during the period preceding the opening of the Gotthard Base Tunnel at the end of 2016 (and, therefore, including 2016). Therefore, we are able to mimic the counterfactual outcome of how the vehicles on the Gotthard motorway section would have evolved in the absence of the Gotthard Base tunnel. The synthetic control method assigns a weight of 16.4% to the Bernina, Flüela, Fréjus Tunnel, Grand-Saint-Bernard, Julier, and Mont-Blanc Tunnel as well as a weight of 1.3% to the San Bernardino Tunnel.
@@
Moreover, Figure 4 also indicates that the number of vehicles (i.e., the average residuals compared to the pre-treatment period) on the Gotthard motorway section were lower compared to the synthetic Gotthard motorway section. On average, the daily number of vehicles was 164 vehicles lower during the months April to October. In other words, the traffic decreased by 1% compared to the pre-treatment average of 16,365 daily vehicles (see Table 1).
To determine the statistical significance of the treatment effect, we randomly draw seven unists with replacement from the donor pool 200 times to arrive at bootstrap confidence intervals (see, e.g., Wallimann, Blättler, and von Arx, 2023, for an example in the transportation literature). The corresponding 95% bootstrap confidence interval of the average estimated effect is [-265; -89] and, thus, does not include zero.
@@ 6.2 Sensitivity analysis
Assumption 4 states that there are no spillover effects from the Gotthard Base Tunnel on other alpine crossings. However, it is conceivable that, for example, a traveler journeying from Hamburg to Milan might alter her route if less congestion is expected on the Gotthard route. For instance, on April 30, 2025, at 11 p.m., travel via the Gotthard takes according to Google Maps 11 hours and 57 minutes, while using the San Bernardino Tunnel takes 11 hours and 48 minutes—making the two routes nearly identical in terms of travel time. Nevertheless, during peak hours, the Gotthard motorway section may now be somewhat less congested, meaning that travelers, after the tunnel’s introduction, may opt not to switch to the San Bernardino Tunnel. To address this potential spillover effect, we exclude the San Bernardino route in our initial sensitivity analysis.
The effect of the new Gotthard Base Tunnel on car traffic, when applying the synthetic control method, amounts to a reduction of 253 vehicles per day, or 1.5% compared to the pre-treatment mean of vehicles, which slightly higher negative than our original result. See also Figure 5 for the development of the average residuals for the Gotthard route and its synthetic counterpart. The 95% bootstrap confidence interval is [–266, -107] and is again statistically significant at the conventional 5% level.
Figure 5: Demand development of the Gotthard motorway section and the synthetic counterpart excluding the San Bernardino
According to Assumption 2, we assume that the units in the donor pool are sufficiently similar to the Gotthard motorway section. To examine whether the introduction of the Gotthard Base Tunnel had an effect on domestic leisure traffic by car, we conducted a second sensitivity analysis by assembling a different donor pool. For this purpose, we selected touristic routes in Switzerland. [^8] Unfortunately, for many routes, data are not available for the full period. Considering yearly averages, the data set contains the Gotthard motorway section and selected counting stations Alpnachstad, Erstfeld, Flüela, Gondo, Gstaad, Hinterrhein, Schattdorf, and Tamis.
In Figure 6 we see that the trajectories of the Gotthard and synthetic Gotthard motorway sections again align closely prior to the intervention. We observe a negative effect of –610 daily vehicles attributable to the Gotthard Base Tunnel. This effect is now noticeably larger negative, corresponding to approximately a 5% reduction in traffic compared to the 2016 average of 13,200 daily vehicles. However, the 95% bootstrap confidence interval for the estimated effect of [-787; 2081] includes zero and the effect should therefore be interpreted with caution.
@@ 7 Discussion and Conclusion
In this study, we employed the synthetic control and synthetic difference-in-differences methods to estimate the causal impact of the Gotthard Base Tunnel the parallel Gotthard motorway section. By constructing a counterfactual based on a weighted combination of alternative alpine crossings in Switzerland and neighboring countries, we estimated that the number of light vehicles decreased by 164 and 152 per day, respectively, during the months from April to October following the tunnel’s opening. Considering the average leisure car occupancy rate of 1.89, this corresponds to approximately 300 individuals per day. These figures represented an approximate 1% decline in vehicles compared to the pre-intervention baseline. Although the magnitude of the effect was modest, bootstrap confidence intervals indicated statistical significance. Further sensitivity analysis accounting for potential spillover effects reinforced these findings. Additionally, when benchmarked against other motorway sections with similar exposure to leisure traffic, the estimated reduction in daily vehicles rose to 610, although not being statistically significant any more. In summary, in contrast to the study of Borsati and Albalate (2020) analyzing the case of high-speed rails (HSR) in Italy, the construction of an efficient railway infrastructure, i.e., the Gotthard Base Tunnel with its travel time savings had a modal shift effect.
@@
Looking at Table 2 in Appendix A, the train remains slower than the car on several routes. Considering access time to the train would make this difference even more pronounced. If the dataset could be restricted to city trips only (or at least to trips originating in urban areas), the estimated effect would likely be relatively larger. Moreover, tourists are often traveling in groups. In such cases, traveling by car is often more cost-effective, especially when none of the passengers hold public transport subscription (e.g., a Half Fare Travelcard). Additionally, with regard to transit travel, there are no direct train connections between Germany and Italy, or vice versa. Therefore, our result is not completely surprising (Chang and Lee, 2008). Also that tourists switch to rail during periods of high north-south (tourist) traffic in order to avoid congestion—thereby reducing peak demand at the road tunnel—cannot be confirmed based on the diurnal traffic patterns shown in Figure B.2 and Figure B.3 in Appendix B.
This short-term perspective constitutes the main limitation of our study. It would be of particular interest to examine the long-term effects, as well as to assess whether the negative impact on the number of vehicles on the parallel Gotthard motorway section intensified following the completion of the Ceneri Base Tunnel, which further increased train frequency and reduced travel times—developments that are also especially relevant for transit traffic, a key segment of travel along the Gotthard corridor. However, this extension of our study was not feasible due to the external shock of the COVID-19 pandemic. Future empirical research on similar natural experiments should aim to investigate the long-term effects of public transport quality improvements on road traffic.
@@ Appendix A Additional Tables
  • Note: Train travel durations are calculated from departure to arrival times. Train connections were derived from the timetable archive available at https://www.oev-info.ch/de/fahrplan-aktuell/fahrplanarchiv. Time savings indicate reductions between 2016 and 2019 but may vary depending on the specific connection. For train journeys, the main station was used as the reference point; for car trips, the location specified by Google Maps was used. Car travel times refer to the shortest suggested driving time according to Google Maps for Monday, April 8, 2019, with a departure at 11:00 AM.

### v2

@@ 6.1 Main Results
Figure 4 shows the main result when applying the synthetic control method. The trajectories of the Gotthard motorway section and the synthetic Gotthard motorway section mainly track each other closely during the period preceding the opening of the Gotthard Base Tunnel at the end of 2016 (and, therefore, including 2016). Therefore, we are able to mimic the counterfactual outcome of how the vehicles on the Gotthard motorway section would have evolved in the absence of the Gotthard Base tunnel. The synthetic control method assigns a non-zero weight of 76.8% to the Bernina and a weight of 23.2% to the Fréjus Tunnel. All other alpine crossings get a weight of 0%.
@@
Moreover, Figure 4 also indicates that the number of vehicles (i.e., the average residuals compared to the pre-treatment period) on the Gotthard motorway section were lower compared to the synthetic Gotthard motorway section. On average, the daily number of vehicles was 135 vehicles lower during the months April to October. In other words, the traffic decreased by 0.8% compared to the pre-treatment average of 16,365 daily vehicles (see Table 1).
To determine the statistical significance of the treatment effect, we randomly draw seven unists with replacement from the donor pool 200 times to arrive at bootstrap confidence intervals (see, e.g., Wallimann, Blättler, and von Arx, 2023, for an example in the transportation literature). The corresponding 95% bootstrap confidence interval of the average estimated effect is [-269; -127] and, thus, does not include zero. Therefore, we conclude that the effect is statistically significantly different from zero at the conventional 5% level.
@@ 6.2 Sensitivity analysis
Assumption 4 states that there are no spillover effects from the Gotthard Base Tunnel on other alpine crossings. To address a potential spillover effect as discussed in Section 4.2, we exclude the San Bernardino route in our first sensitivity analysis.
The effect of the new Gotthard Base Tunnel on car traffic, when applying the synthetic control method, amounts to a reduction of 135 vehicles per day, or 0.8% compared to the pre-treatment mean of vehicles, which is the same as in the original result, as the San Bernardino got a zero weight. The 95% bootstrap confidence interval now amounts to [–269, -110] and is again statistically significant at the conventional 5% level.
Note that when using the synthetic control method to construct our Synthetic Gotthard motorway section, we include the pre-treatment values of the dependent variable for all years in the pre-treatment period from 2013 to 2016. This allows the method to match the entire pre-treatment trajectory, rather than just the average. Therefore, as a second sensitivity analysis, we exclude the pre-treatment predictors and instead match solely on the outcome variable over the full pre-treatment period. [^9] This corresponds to deactivating the special.predictors argument in the dataprep command in the statistical software R. The estimated effect of the new Gotthard Base Tunnel on traffic amounts now to $-152$ vehicles per day—or approximately 1% relative to pre-treatment traffic levels—with a 95% bootstrap confidence interval of [$-265$; $-89$].
According to Assumption 2, we assume that the units in the donor pool are sufficiently similar to the Gotthard motorway section. To examine whether the introduction of the Gotthard Base Tunnel had an effect on domestic leisure traffic by car, we conducted a third sensitivity analysis by assembling a different donor pool. For this purpose, we selected touristic routes in Switzerland. [^10] Unfortunately, for many routes, data are not available for the full period. Considering yearly averages, the data set contains the Gotthard motorway section and selected counting stations Alpnachstad, Erstfeld, Flüela, Gondo, Gstaad, Hinterrhein, Schattdorf, and Tamis.
In Figure 5 we see that the trajectories of the Gotthard and synthetic Gotthard motorway sections again align closely prior to the intervention. We observe a negative effect of –617 daily vehicles attributable to the Gotthard Base Tunnel. This effect is now noticeably larger negative, corresponding to approximately a 5% reduction in traffic compared to the 2016 average of 13,200 daily vehicles. However, the 95% bootstrap confidence interval for the estimated effect of [-785; 2008] includes zero and the effect should therefore be interpreted with caution.
@@ 7 Discussion and Conclusion
In this study, we employed the synthetic control and synthetic difference-in-differences methods to estimate the causal impact of the Gotthard Base Tunnel on the parallel Gotthard motorway section. By constructing a counterfactual based on a weighted combination of alternative alpine crossings in Switzerland and neighboring countries, we estimated that the number of light vehicles decreased by 135 and 152 per day, respectively, during the months from April to October following the tunnel’s opening. Considering the average leisure car occupancy rate of 1.89, this corresponds to approximately 270 individuals per day. These figures represented an a decline in vehicles of just under 1% compared to the pre-intervention baseline. Although the magnitude of the effect was modest, bootstrap confidence intervals indicated statistical significance. Further sensitivity analysis, e.g., accounting for potential spillover effects reinforced these findings. Additionally, when constructing a counterfactual with other Swiss motorway sections with similar exposure to leisure traffic—instead of using other alpine crossings in Switzerland and neighboring countries—, the estimated reduction in daily vehicles rose to 617. However, it is no longer statistically significant. In summary, in contrast to the study of Borsati and Albalate (2020) analyzing the case of high-speed rails (HSR) in Italy, the construction of an efficient railway infrastructure, i.e., the Gotthard Base Tunnel with its travel time savings had a modal shift effect.
@@
However, with the Gotthard Base Tunnel, only pull measures were implemented. When looking at mode share on the Gotthard route, we observe a rise in the public transport share from 25.7% in 2016 to 31.4% in 2019 (see Figure 6). However, only approximately 270 individuals shifted from car to public transport per day, corresponding to just 2.3% of all public transport users. Moreover, assuming identical growth rates as in the Swiss long-distance public transport sector, 80.4% of users would have traveled by public transport anyway. This leaves 17.3% as induced demand, i.e., new public transport users who would not have taken the Gotthard route in the absence of the Gotthard Base Tunnel. In conclusion, achieving a substantial modal shift through pull measures alone is difficult; they must be complemented by push measures such as road pricing.
Figure 6: Decomposition of the 2019 public transport share
Looking at Table 2 in Appendix A, the train remains slower than the car on several routes. Therefore, if the dataset could be restricted to city trips only (or at least to trips originating in urban areas), the estimated effect would likely be relatively larger. Moreover, tourists are often traveling in groups. In such cases, traveling by car is often more cost-effective, especially when none of the passengers hold public transport subscription (e.g., a Half Fare Travelcard). Additionally, with regard to transit travel, there are no direct train connections between Germany and Italy, or vice versa. Therefore, our result is not completely surprising (Chang and Lee, 2008). Also that tourists switch to rail during periods of high north-south (tourist) traffic in order to avoid congestion—thereby reducing peak demand at the road tunnel—cannot be confirmed based on the diurnal traffic patterns shown in Figure B.2 and Figure B.3 in Appendix B.
The short-term perspective, spanning three years, outlined in this paper constitutes the main limitation of our study. It would be of particular interest to examine the long-term effects (e.g., more than ten years), as well as to assess whether the negative impact on the number of vehicles on the parallel Gotthard motorway section intensified following the completion of the Ceneri Base Tunnel, which further increased train frequency and reduced travel times—developments that are also especially relevant for transit traffic, a key segment of travel along the Gotthard corridor. However, this extension of our study was not feasible due to the external shock of the COVID-19 pandemic. Future empirical research on similar natural experiments should aim to investigate the long-term effects of public transport quality improvements on road traffic.
@@ Appendix A Additional Tables
  • Note: Train travel durations are calculated from departure to arrival times. Train connections were derived from the timetable archive available at https://www.oev-info.ch/de/fahrplan-aktuell/fahrplanarchiv. Time savings indicate reductions between 2016 and 2019 but may vary depending on the specific connection. For train journeys, the main station was used as the reference point; for car trips, the location specified by Google Maps was used. Car travel times refer to the shortest suggested driving time according to Google Maps for Monday, April 8, 2019, with a departure at 11:00 AM. Note that if we would consider access time to the train would make this difference even more pronounced.